# Exploiting the Uncoordinated Privacy Protections of Eye Tracking and VR Motion Data for Unauthorized User Identification

### v1

@@ 4.2.1 Gaze Data Privacy Mechanism
We employ a privacy mechanism for gaze data proposed by Wilson et al. [33], which obfuscates personally identifying information by applying a smoothing operation to the gaze data as it is streamed from the device. To smooth the data, each gaze sample is re-defined as the linearly weighted average of $B$ preceding samples. For this investigation, we set the parameter $B=108$, which corresponds to the 1.2 seconds of data that the original work was validated on.
@@ 4.2.2 Head and Hand Data Privacy Mechanism
We employ a portion of MetaGuard from Nair et al. [23] to serve as our privacy mechanism for our headset and hand data. MetaGuard applies $\epsilon$-differential privacy to obscure physical characteristics of the user’s body, thereby reducing the chance that users are re-identified based on their physical traits. We utilize the portion of their method that obfuscates the users’ height and wingspan by adding bounded Laplacian noise to the vertical component of headset data and to the hand data, respectively. To estimate the upper and lower bounds for height and wingspan in our population, we approximate each user’s height based on OpenNEEDS headset data. Specifically, we define each user’s height as the average vertical component of their headset position and assume this value also represents their wingspan, given a typical 1:1 height-to-wingspan ratio. This approach provides a reasonable way to approximate the height and wingspan bounds needed to constrain privatized data; from this approach, we constrain the heights and wingspans in our privatized data to a range of [1.32, 1.82] meters.
We do not perturb the horizontal or depth components of our headset data, because the original method does not do so unless it is to protect an adversary from inferring the dimensions of the user’s play area—because all of our users utilized the same space and the justification for perturbing these components has little relevance for personal identity, we determined that perturbing these data streams would represent an improper deviation from the original method. For this investigation, we enforce parity on the privatization statues of VR motion data—in other words, if head and hand data are both available, they are either both privatized, or both unprivatized.
@@ 4.3 Biometric Authentication Model
We employ EyeKnowYouToo (EKYT) [17], a model for eye movement biometrics, as our biometric authentication model. While EKYT was originally developed and validated for gaze-based biometrics, we extend EKYT to perform biometric authentication using headset and hand motion data by adding them as additional channels to the input.
To test the effect of combining data streams on biometric performance, we train and evaluate 7 models on unprivatized (unmodified) data. These models establish a baseline of biometric performance for this investigation, and their results are evaluated in Section 5.1 (Experiment IDs 01-07). We also train and evaluate an additional 7 models on “fully” privatized data (i.e., where each data stream has been processed by a privacy mechanism) to establish identification rates under the best-case privacy protection that we can expect from this investigation. These models are evaluated in Section 5.2 (Experiment IDs 08-14).
To test the ability of the biometric model to circumvent partial privacy protections with other sources of data, we train and test an additional 6 models on various combinations of privatized and unprivatized data steams, where at least one data stream is not privatized. These models investigate the extent to which EKYT can leverage alternative sources of data to circumvent privacy protections, and are evaluated in Section 5.3 (Experiment IDs 15-20).
@@ 4.3.1 Data Pre-Processing
Eye movement data was processed using the methodology described in Lohr and Komogortsev’s original implementation of EKYT [17]. Given the positional gaze data provided by the VR device, we first calculate the velocity of the eye movement using a Savitzky-Golay differentiation filter with an order of 2 and a window size of 7. Each sample then contained the horizontal and vertical velocity of the eye at a point in time, denoted as $G_{t}=(v_{x},v_{y})_{t}$. The velocity data is then split into non-overlapping windows of 5 seconds, which corresponds to 450 samples at 90 Hz. In each window, velocities were clamped to physiologically realistic movement speeds of $\pm$1000 degrees of visual angle per second (/sec), then normalized using the mean and standard deviation of eye movement velocities derived from the training data and replaced NaN values with 0. These processed velocity windows were then provided to EKYT for training and evaluation.
During experimentation, we observed no particular benefit of extensively pre-processing the VR motion data. As a result, we simply split the raw positional data into non-overlapping windows of 5 seconds, and provided this input to the model for training and evaluation. We provided the model with positional headset data that contained the horizontal, vertical, and depth position of the headset in world space, denoted as $H_{t}=(h_{x},h_{y},h_{z})_{t}$. Although we have rotational data available from the headset as well, we chose to only provide headset positions as input in order to prevent information from any single data stream from being overrepresented in the inputs. Hand data was provided in the form $(L,R)_{t}=(\{l_{x},l_{y},l_{z}\},\{r_{x},r_{y},r_{z}\})_{t}$, which respectively denote the horizontal, vertical, and depth of the left and right hands in world space at time $t$. We only provide positional data of the hands to the model as input, but we note that including rotational data and interaction patterns (e.g., button presses, gestures) may also contain uniquely identifiable information [27].
@@ 4.3.2 Training and Evaluation
We train and test separate authentication models for each test condition described in Section 4.3. Each model was trained for up to 100 epochs, with a learning rate of 3e-4. Due to the limited number of subjects in our dataset, each model is trained using four-fold cross-validation. Each fold is trained on 75% of subjects and evaluated on the remaining 25% of subjects that were held out, ensuring that the training and evaluation sets are subject-disjoint. We report each model’s biometric performance as the mean authentication performance on the held-out subjects across all four folds.
We report biometric authentication performance in two settings—biometric verification and biometric identification. During biometric verification, EKYT receives both a sample of data and a claimed user identity, and compares that sample to the user’s enrollment template to determine whether the sample could have been generated by the claimed identity. The system then returns a Yes/No decision if the sample is sufficiently similar to the enrollment template. We measure biometric verification performance using Equal Error Rate (EER). Generally, lower EER indicates better biometric performance. During biometric identification, EKYT receives a sample of data and returns the identity of the user who most likely generated the data. We report biometric identification performance as the Rank-1 Identification Rate (Rank-1 IR), or the percentage of times that EKYT returns the correct identity. Generally, higher Rank-1 IR indicates better biometric performance.
We first establish baseline biometric performance when using combinations of unmodified gaze, head, and hand data, then then observe how that biometric performance changes after applying privacy mechanisms to each data stream. We then assess a “best-case” scenario where all three data streams are privatized, and observe the extent to which privacy may be preserved using relevant state-of-the-art privacy-enhancing mechanisms. In an ideal scenario, a biometric model trained on a privatized data stream would produce chance-level biometric performance at evaluation time, which corresponds to 50% EER for verification and $1/38\approx 2.6\%$ IR for identification. Finally, we explore how the privacy gains afforded by these mechanisms may be circumvented by only applying privacy mechanisms to a subset of available data, and allowing EKYT access to all other unmodified data streams. This evaluation configuration represents a scenario that may be encountered in real-world VR systems, where a patchwork of privacy protections are applied without consideration for other sources of risk.
@@ 5.1 Combining Unmodified Data Streams for User Authentication
We first establish a baseline for EKYT’s biometric authentication performance when it is trained and tested on unmodified data from OpenNEEDS. These results, which are summarized in Table 1, can represent the risk of user identification in a scenario where a VR platform streams sensor data to third party applications without applying additional privacy protection.
@@
Firstly, using each of the data streams in isolation produces authentication rates that are well above chance-level performance, which indicates the EKYT can meaningfully extract user-specific information to re-identify them. On average, gaze data (E01) and head positions (E02) produce similar authentication performance when used in isolation. On the other hand, it appears that hand positions (E03) were not as informative of user identity; we observe a notable decrease in biometric performance across both authentication settings relative to headset and gaze data.
@@
Combining two or more data sources does not appear to improve biometric performance over utilizing any single data stream, but it does help overcome poor performance when a more informative data stream (e.g., gaze or head) is supplied to the model alongside a less informative data stream (e.g., hands). Because gaze-only and head-only biometric performance produced such similar results in isolation, it is perhaps unsurprising that they do not produce similar biometric performance when combined—we observe a slight (but likely not significant) deterioration in authentication performance between E01/E02 and E05 after combining gaze and head data, and observe no significant change in Rank-1 IR.
@@
However, combining either gaze or head data with hand data can improve the biometric performance over using hand data alone. Combining gaze with hand data, or head with hand data, produces a notable improvement in biometric performance on average—EER decreases from 30% (E03) to as low as 19.2% (E04) on average, and identification rate increases from 42.5% to as high as 65%. Finally, combining all three data streams (E07) does not appear to significantly affect verification rates over the best individual data stream, but does improve the identification rate to its highest observed value of 75%. The improvement in identification rate suggest that combining data streams may give the biometric model more opportunity to identify unique behavior across all three streams compared to using a single data stream in isolation, as person-specific idiosyncrasies would be captured simultaneously in all three sources of data. These findings indicate that combining data streams that are known to convey identifying information may not necessarily harm biometric performance, even when combining richer data streams (such as gaze and head data) with less informative data (e.g., hand data).
@@ 5.2 User Authentication with Fully Privatized Data
Next, we explore the impact of applying the privacy mechanisms from Wilson et al. [33] and Nair et al. [23] to gaze and VR motion data, respectively. When applied to their respective data streams, these privacy mechanisms obfuscate the personal information that is otherwise encoded in the data. The resulting “privatized” data is ideally absent of personally identifying features. This renders EKYT unable to learn meaningful user-specific patterns during training, thereby reducing the chances that users would be recognized across sessions. For ideal privacy protection (i.e., where no user can be re-identified), the resulting biometric authentication performance would be at or near chance-level performance, which corresponds to 50% EER and 2.6% IR for verification and identification, respectively. In this section, we explore a best-case interaction scenario where all of the potentially sensitive data streams have undergone privacy enhancement before being provided to an application. Table 2 shows the biometric performance of EKYT models that were trained on this privatized data.
@@
First, we evaluate the effectiveness of each privacy mechanism at reducing authentication rates compared to unmodified data (visualized in Figure 2). The gaze-based privacy mechanism (E08) is the most effective at obscuring personal information from EKYT, because it degrades biometric performance closest to chance-level verification rate (47.2% EER) and chance-level identification (10% IR) compared to its unmodified counterpart, E01 (EER=18.9%, IR=62.5%). The privacy mechanism for head data produces a relatively more modest level of privacy enhancement, degrading authentication performance from 18.6% EER and 60% IR (E02) to 32.5% EER and 43.9% IR (E09). This degradation in authentication performance indicates that EKYT is less capable of recognizing users across sessions, which represents an increase in privacy relative to the unmodified data. The privacy mechanism applied to the hands is less effective at reducing authentication rates, and therefore produce comparable (or slightly better) biometric performance relative to unmodified data (E03 and E10).
@@
Overall, the biometric performance of the combined privatized data streams (E04-E07) is notably lower than their unmodified counterpart (E11-E14), which indicates some privacy enhancement is afforded when all of the available data streams undergo privacy enhancement. However, authentication performance for these combined data sources tends to reproduce the performance of the least effectively privatized data stream. For example, when provided with privatized gaze and head data (E12), EKYT produces comparable biometric performance to E09, where only privatized head data is available. This can also be observed when comparing performance after supplying EKYT with privatized gaze and hand data (E13), solely gaze data (E08), and solely hand data (E10). While the increased privacy protection for gaze data prevents EKYT from learning user-specific information from the user’s gaze, it did not degrade biometric performance when EKYT has access to other, less effectively privatized data streams, as user-specific information was still learnable from these alternative sources. From these results, it is reasonable to consider that comprehensive privacy defenses are only as effective as their “weakest link,” or least effective component.
@@
Figure 2: A visualization summarizing the effect of privatization on biometric authentication performance in both the verification (left) and identification (right) settings.
5.3 Exploiting Partial Privatization to Circumvent Privacy Protections
@@
After establishing the upper and lower bound of user authentication rates for our data set (Sections 5.1 and 5.2, respectively), we now investigate the privacy landscape between these extremes. In real-world VR applications, it cannot be assumed that comprehensive privacy protections have been put into place for every application. While the worst-case scenario is that there are no privacy protections in place, we are more likely to encounter inconsistently applied privacy protections across applications, where privacy protections have been applied to some potentially identifying data streams, but not to others. The experiments in this section represent this patchwork approach to privacy enhancement, where some potentially sensitive data streams have undergone privacy enhancement prior to being shared and others are provided to third party applications without any additional protections applied. We are interested in whether a biometric authentication model can leverage these unprivatized data streams to circumvent the privacy enhancements afforded by the privatized data, thereby defeating the purpose of privatization altogether. If this is the case, then the model would produce user authentication rates that are better the high-privacy scenario (E08-E14), and potentially reach comparable performance levels as the no-privacy scenario (E01-E07).
@@
Table 3 shows the biometric performance of EKYT that was trained on various combinations of privatized and non-privatized data. Figure 3 visually summarizes the difference in performance for related test conditions. Recall that, when head data and hand data are both available, they are either both privatized or both unmodified.
@@
We previously observed that privatized gaze data produced the most privacy enhancement (E08), while the privatized head data produced relatively more modest but still demonstrable levels of privacy protection. When providing privatized gaze data and unmodified head data (E15), EKYT produces biometric authentication rates closest to E05, where no privacy mechanism has been applied at all. This indicates that the privacy gains afforded by the privatized gaze data were significantly impacted by the presence of unmodified head data, and vice versa. We also observe this effect, albeit to a lesser extent, when EKYT is provided privatized head data and unmodified gaze data (E18, Figure 3(b)). When EKYT is unable to identify users through their gaze data, it is able to instead leverage information from the headset to successfully perform biometric authentication. While headset data (E15) appears to be more effective at circumventing privatized gaze data than hand data (E16), combining both head and hand data (E17) also produces notably high-quality biometric authentication results (though we would attribute this to the headset data making the most contribution to biometric performance). Figure 3(a) visualizes the extent to which headset and hand data improve biometric performance when combined with privatized gaze data.
We observed that the privacy mechanisms that were applied to hand data were not particularly effective at obscuring user identities from EKYT—biometric performance did not change significantly between unmodified and privatized hand data. Our attempts to circumvent privatized hand data using gaze (E19) data produced performance values close to the corresponding baseline experiments (E06), but attempting to circumvent privacy using all data streams (E20) resulted in significantly degraded performance relative to the baseline experiment (E07), likely due to the additional privatization of the head data. Figure 3(c) visualizes the extent to which headset and hand data improve biometric performance when combined with privatized gaze data.
@@
(a) Exploiting unmodified head and hand data to circumvent privacy protections for gaze.
(b) Exploiting unmodified gaze data to circumvent privacy protections for head data. Note that, when headset data is privatized, hand data is also privatized.
(c) Exploiting unmodified gaze data to circumvent privacy protections for hand data. Note that, when hand data is privatized, headset data is also privatized.
Figure 3: A series of bar charts visualizing the degree to which each privatized data stream can be circumvented with different combinations of unmodified data. The most effective circumvention of privacy protections is represented by a low EER score and a high IR score; the most effective configuration for bypassing each privatized data stream has been marked with a star.

### v3

@@ 4.2.1 Gaze Data Privacy Mechanism
We employ privacy mechanism proposed by Wilson et al. [41] for gaze data, which obfuscates personally identifying information in gaze data by applying a smoothing operation to the gaze data as it is streamed from the device. To smooth the data, each gaze sample is re-defined as the linearly weighted average of $B$ preceding samples. For this investigation, we set the parameter $B=108$, which corresponds to the 1.2 seconds of data that the original work was validated on.
@@ 4.2.2 Head and Hand Data Privacy Mechanism
We employ a portion of MetaGuard from Nair et al. [31] to serve as our privacy mechanism for our headset and hand data. MetaGuard applies $\epsilon$-differential privacy to obscure physical characteristics of the user’s body, thereby reducing the chance that users are re-identified based on their physical traits. We utilize the portion of their method that obfuscates the users’ height and wingspan by adding bounded Laplacian noise to the vertical component of headset data and to the positional hand data, respectively. We estimate the upper and lower bounds of the noise distributions by approximating the height of the users in OpenNEEDS. To do so, we designate the height and wingspan of each user as the average vertical component of their headset data during their first recording, assuming a 1:1 height-to-wingspan ratio. Using this method, we then constrain the range of both heights and wingspans in our data to [1.32, 1.82] meters and apply bounded Laplacian noise with these bounds. Following the privacy budgets proposed in the original work, we set $\epsilon=1$ for headset data and $\epsilon=0.5$ for hand data.
We do not perturb the horizontal or depth components of our headset data, because the original method only does so to prevent an adversary from inferring the dimensions of the user’s play area—because all of our users utilized the same space and the justification for perturbing these components has little relevance for personal identity, we determined that perturbing these data streams would not be necessary for our problem setting.
For this investigation, we enforce parity on the privatization statues of VR motion data—in other words, if head and hand data are both available, they are either both privatized, or both unprivatized.
@@ 4.3 Biometric Authentication Model
We employ EyeKnowYouToo (EKYT) [22], the state-of-the-art biometric model for eye movement biometrics, as our biometric authentication model. While EKYT was originally developed and validated for gaze-based biometrics, we extend EKYT to perform biometric authentication using headset and hand motion data by adding them as additional channels to the input.
To test the effect of combining data streams on biometric performance, we train and evaluate 7 models on unprivatized (unmodified) data. These models establish a baseline of biometric performance for this investigation, and their results are evaluated in Section 5.1 (Experiment IDs 01-07). We train and evaluate an additional 7 models on fully privatized data to establish identification rates under the best-case privacy protection that we can expect from this investigation. In this context, this means that every available data stream has been processed by its respective privacy mechanism. These models are evaluated in Section 5.2 (Experiment IDs 08-14).
To test the ability of the biometric model to circumvent incomplete privacy protections with other sources of data, we train and test an additional 6 models on various combinations of privatized and unprivatized data steams, where at least one data stream is not privatized. These models investigate the extent to which EKYT can leverage alternative sources of data to circumvent privacy protections, and are evaluated in Section 5.3 (Experiment IDs 15-20).
@@ 4.3.1 Data Pre-Processing
Eye movement data was processed according the methodology described in EKYT’s original implementation [22]. Given the positional gaze data provided by the VR device, we first calculate the velocity of the eye movement using a Savitzky-Golay differentiation filter with an order of 2 and a window size of 7. The velocity data is then split into non-overlapping windows of 5 seconds, which corresponds to 450 samples at 90 Hz. In each window, these velocities were clamped to realistic movement speeds of $\pm$1000/sec, then normalized using the mean and standard deviation of eye movement velocities derived from the training data and replaced NaN values with 0. These processed velocity windows were then provided to EKYT for training and evaluation. As a result, EKYT is provided windows of 450 samples, where each sample contains the horizontal and vertical velocity of the eye at a point in time, denoted as $G_{t}=(v_{x},v_{y})_{t}$.
During experimentation, we observed no particular benefit of extensively pre-processing VR motion data. VR motion data was split into non-overlapping windows of 5 seconds and provided to the model for training and evaluation. The model received 450-sample windows of positional headset data, where each sample contained the horizontal, vertical, and depth position of the headset in world space, denoted as $H_{t}=(h_{x},h_{y},h_{z})_{t}$. Although we have rotational data available from the headset as well, we chose to only provide headset positions as input in order to prevent information from any single data stream from being overrepresented in the inputs. Hand data was also provided in 450-sample windows, with each sample containing $(L,R)_{t}=(\{l_{x},l_{y},l_{z}\},\{r_{x},r_{y},r_{z}\})_{t}$, which respectively denote the horizontal, vertical, and depth of the left and right hands in world space at time $t$. We only provide positional data of the hands to the model as input, but we acknowledge that including rotational data and interaction patterns (e.g., button presses, gestures) may also contain information that can uniquely identify a user [35].
@@ 4.3.2 Training and Evaluation
We train and test separate authentication models for each test condition described in Section 4.3. Each model was trained for 100 epochs, with a learning rate of 3e-4. Due to the limited number of subjects in our dataset, each model is trained using four-fold cross-validation. Each fold is trained on 75% of subjects and evaluated on the remaining 25% of subjects that were held out, ensuring that the training and evaluation sets are subject-disjoint. We report each model’s biometric performance as the mean authentication performance on the held-out subjects across all four folds.
We evaluate biometric authentication performance in two settings—biometric verification and biometric identification. During biometric verification, EKYT receives both a sample of data and a claimed user identity, and compares that sample to the user’s enrollment template to determine whether the sample could have been generated by the claimed identity. The system then decides whether the sample is sufficiently similar to the enrollment template, returning “True” if the samples match and “False” otherwise. We summarize biometric verification performance using Equal Error Rate (EER). Generally, lower EER indicates better biometric performance. During biometric identification, EKYT receives a sample of data and returns the identity of the user who most likely generated the data. We report biometric identification performance as the Rank-1 Identification Rate (Rank-1 IR), or the percentage of times that EKYT returns the correct identity. Generally, higher Rank-1 IR indicates better biometric performance.
In an ideal scenario, data that has undergone privatization would not contain any identifying information. A biometric model would therefore produce chance-level biometric performance when using this data, which corresponds to 50% EER for verification and $1/38\approx 2.6\%$ IR for identification. To assess the effect of privatization on our data set, we first establish baseline biometric performance when using combinations of unmodified gaze, head, and hand data, then then observe how that biometric performance changes after applying privacy mechanisms to different data streams. We then assess how the privacy gains afforded by these mechanisms may be circumvented by only applying privacy mechanisms to a subset of available data, and allowing EKYT access to all other unmodified data streams.
@@ 5.1 Combining Unmodified Data Streams for User Authentication
We first establish a baseline for EKYT’s biometric authentication performance when it is trained and tested on unmodified data from OpenNEEDS. The authentication rates observed from unmodified data, summarized in Table 1 and Figures 2 and 3, represents the risk of user identification in a scenario where a VR platform streams sensor data to third party applications without applying privacy protections. We present biometric performance when EKYT has access to a single data stream (E01-E03), two data streams (E04-E06), and all three data streams (E07) to explore the effect of combining unmodified data streams on user authentication rates.
@@
We first observe that providing a single data stream (E01-E03) produces authentication rates that are well above chance-level performance, which indicates the EKYT can meaningfully extract user-specific information that can be used to re-identify them across play sessions. Gaze data (E01) and headset position data (E02) produce the similar rates of user authentication, and yield similar verification and identification rates on average. On the other hand, it appears that hand positions (E03) were not as informative for inferring user identity; we observe a notable decrease in biometric performance across both authentication settings relative to gaze and headset data. While features that are related to hand motion data have shown promise for use as a soft biometric (e.g., hand dominance and derivable body proportions [35]), there is relatively less evidence that raw hand motion data alone is as suitable for user identification as gaze data [22, 23] and headset data [18, 40].
@@
We next observe the effect of combining two sources of data on user authentication rates. Overall, combining two data sources does not appear to improve biometric performance over utilizing any single data stream, but it can help overcome poor biometric performance when a more informative data stream (e.g., gaze or head) is supplied to the model alongside a less informative data stream (e.g., hands). For example, combining head and hand data (E04, known hereafter as “VR motion data”) on authentication rates yields biometric authentication performance similar comparable to the test condition where only headset data is available (E02), which is a notable improvement in performance relative to the hands-only condition (E03). Similarly, combining gaze and hand data produces a negligible degradation in biometric performance on average relative to the gaze-only condition (E06 versus E01), indicating that EKYT can effectively leverage the more informative data stream to meaningfully distinguish between users. We also observe a slight (but likely not significant) deterioration in authentication performance between E01/E02 and E05 after combining gaze and head data, indicating that EKYT did not distinguish users more meaningfully on average when these data sources were combined.
@@
Finally, combining all three data streams (E07) does not appear to significantly affect verification rates over the best individual data stream, but does improve the identification rate to its highest observed value of 75%. The modest improvement in identification rate suggest that combining data streams may give the biometric model more opportunity to identify unique behavior across all three streams compared to using a single data stream in isolation, as person-specific idiosyncrasies would be captured simultaneously in all three sources of data.
@@ 5.2 User Authentication with Fully Privatized Data
We next explore a best-case interaction scenario where all of the data streams that we consider have undergone privacy enhancement. Specifically, we explore the impact of applying Wilson et al. [41]’s smoothing mechanism to the gaze data and applying Nair et al. [31]’s MetaGuard mechanism to VR motion data, respectively. When applied to their respective data streams, these privacy mechanisms obfuscate the personal information encoded in the data, producing “privatized” data is ideally absent of personally identifying features. This hinders EKYT’s ability to learn meaningful user-specific patterns during training, thereby reducing the chances that users can be re-identified across play sessions. In an ideal scenario where the privatized data is totally absent of personally identifying information, the resulting biometric authentication rates would be at or near chance-level performance, which corresponds to 50% EER and 2.6% IR for verification and identification, respectively. We observe the impact of privatization on degrading EKYT’s biometric authentication rates toward this ideal. Table 2 shows the average biometric performance of EKYT models that were trained on privatized data, while Figures 2 and 3 show biometric performance for identification and verification, respectively.
@@
First, we evaluate the effectiveness of each privacy mechanism at reducing authentication rates for each data stream individually. The gaze-based privacy mechanism (E08) is the most effective at obscuring personal information from EKYT, degrading biometric performance closest to chance-level verification rates (47.2% EER versus 50% at chance) and identification rates (10% IR versus 2.6% at chance), as well as producing the greatest reduction in authentication rates relative to its unmodified counterpart (E01). The privacy mechanism for head data produces relatively more modest privacy enhancement, degrading authentication performance from 18.6% EER and 60% IR (E02) to 31.7% EER and 38.9% IR (E09). This degradation in authentication performance indicates that EKYT is less capable of recognizing users across sessions, which represents an increase in privacy relative to the unmodified data. On the other hand, the privacy mechanism applied to the hand data is less effective at reducing authentication rates, and therefore produces comparable (or even slightly better) biometric performance relative to unmodified data (E03 versus E10). We theorize that the limited privacy protections we observe for our hand data stems from unique characteristics of both EKYT and our privacy mechanism of choice; while MetaGuard is designed obscures the physical properties that can be derived from a user’s hand data, such as their wingspan and arm length [31], it does not obscure motion patterns that can be used for identification on a short timescale [27]. Because EKYT learns to extract such behavioral patterns by leveraging minute changes in gaze velocity trajectory [22], it is possible that EKYT can overcome our implementation of MetaGuard by extracting user-specific information from behavioral patterns rather than from physicality. It it also possible that the privacy budget employed for MetaGuard is not an optimal choice for the OpenNEEDS data set, as it was derived for another data set that utilized a different set of stimuli [31].
@@
Overall, the biometric performance when combining privatized data streams (E11-E14) is notably worse than their unmodified counterparts (E04-E07), which indicates some privacy enhancement is afforded when all of the available data streams undergo privacy enhancement. However, authentication performance for these combined data sources tend to reproduce the performance of the least effectively privatized data stream. For example, when provided with privatized gaze and hand data (E13), EKYT produces comparable biometric performance to E10, where only the poorly privatized hand data is available. While the increased privacy protection for gaze data prevents EKYT from learning user-specific information from the user’s gaze, it did not degrade biometric performance when EKYT has access to other, less effectively privatized data streams, as user-specific information was still learnable from these alternative sources. From these results, it’s reasonable to consider that comprehensive privacy defenses are only as effective as their “weakest link,” or least effectively privatized component.
@@
Figure 2: A visualization summarizing the effect of privatization on biometric identification rates.
(a) Gaze data.
(b) Headset data.
(c) Hand data.
(d) VR motion data.
Figure 3: ROC curves comparing the effect of privatization on biometric verification rates for each data stream individually, and for VR motion data.
5.3 Circumventing Partial Privacy Protections for User Authentication
@@
After establishing the upper and lower bound of user authentication rates for our data set (Sections 5.1 and 5.2, respectively), we now investigate the privacy landscape that exists between these extremes. In real-world VR applications, it cannot be assumed that comprehensive privacy protections are implemented or respected by every application, if these protections were not applied before the data was shared in the first place. While the worst-case scenario is that there are no privacy protections in place, we are more likely to encounter inconsistently applied privacy protections across applications, where privacy protections have been applied to some potentially identifying data streams, but not to others. The experiments in this section represent this patchwork approach to privacy enhancement, where some potentially sensitive data streams have undergone privacy enhancement prior to being shared and others are provided without any additional protections. We are interested in whether EKYT can leverage these unprivatized data streams to circumvent the privacy enhancements afforded by the privatized data, thereby defeating the purpose of privatization altogether. If this is the case, then the model would produce user authentication rates that are better the high-privacy scenario (E08-E14), and potentially reach comparable performance levels as the no-privacy scenario (E01-E07).
@@
Table 3 shows the biometric performance of EKYT that was trained on various combinations of privatized and non-privatized data. Recall that, when head data and hand data are both available, they are either both privatized or both unmodified. Figure 4 visually summarizes the difference in identification rates for related test conditions, and Figure 5 compares verification performance among the same conditions.
@@
We first examine biometric performance when both gaze data and head data are available, but privatization is applied only to gaze data (E15) or head data (E18). Figures 4(a) and 5(a) visualize the extent to which each unmodified data stream achieves this circumvention. We previously observed that unmodified gaze data and head data produce similar authentication rates, while privatized gaze data produced better privacy enhancement than privatized head data. However, when providing privatized gaze data and unmodified head data (E15), EKYT produces biometric authentication rates that are comparable to a condition where no privacy mechanism had been applied at all. We also observe this effect, albeit to a lesser extent, when EKYT is provided privatized head data and unmodified gaze data (E18). Despite the demonstrable levels of privacy enhancement afforded to both data streams individually, EKYT was able to leverage the information encoded in the unmodified data stream to successfully authenticate users.
Next, we observe biometric performance when both gaze data and hand data are available, but privatization is applied only to gaze data (E16) or hand data (E19). Figures 4(b) and 5(b) visualize the extent to which each unmodified data stream impacts user authentication rates. We previously observed that hand data is the least effective data stream for user authentication, and that the privacy mechanism applied to is not particularly effective at obscuring user identities from EKYT—biometric performance did not change significantly between unmodified and privatized hand data. As a result, attempting to bypass privacy protections applied to gaze data using unmodified hand data (E16) does not significantly improve biometric authentication rates over either the fully privatized data (E13) nor the unmodified baseline performance of hand-only data (E03). Because hand data has limited utility as a biometric in this investigation, this attempt to circumvent privacy protections would yield similarly limited levels success. On the other hand, attempting to circumvent the privacy protections applied to hand data using unmodified gaze data (E19) produces significantly better biometric authentication rates than not only the fully privatized data (E13), but also the test condition where neither data stream was privatized (E06). While this result is perhaps unsurprising, it highlights that both poorly implemented privacy protections and the complete absence of privacy enhancement can increase the risk of unauthorized user authentication.
Finally, we observe the biometric performance when both gaze and VR motion data (i.e., head and hand data) are available, but only gaze data (E17) or VR motion data (E20) are privatized. Figures 4(c) and 5(c) visualize the extent to which each unmodified data stream impacts user authentication rates. EKYT can effectively circumvent privacy protections when using unmodified VR motion data (E17), producing similar authentication performance as most other experiments that utilize unmodified headset data. Interestingly, the experiment circumventing privacy protections with unmodified VR motion data (E17) appears to yield significantly better authentication rates than the experiment doing so with unmodified gaze data (E20). Gaze data is significantly less effective at circumventing privacy protections when VR motion data is privatized (E20)—when EKYT only has unmodified gaze data at its disposal, it produces authentication performance at levels that are comparable to a scenario where all three data streams have been privatized (E14). In other words, the unmodified gaze data was not effective for user authentication in the presence of privatized VR motion data. We attribute this finding to the over-representation of VR motion data in EKYT at training time—when training on all three data streams, gaze data only contributes two channels of input and VR motion data collectively contributes nine (see Section 4.3.1 for details). Because nine of the eleven total channels provided to EKYT were privatized and therefore not informative for biometric authentication in E20, EKYT likely disregarded the user-specific behavior that was only represented in a minority of input channels at training time. Conversely, this also explains why unmodified VR motion data is highly effective at circumventing privacy protections in E17—the number of input channels that conveyed user-specific behavior outnumbered the number of privatized input channels.
@@
(a) Gaze and head data.
(b) Gaze and hand data.
(c) Gaze and VR motion data
Figure 4: The impact of incomplete privatization of each combination of data streams on user identification rates (yellow). Blue bars display identification rates when all data has undergone privatization, and red bars display identification rates when all of the data is unmodified.
(a) Gaze and head data.
(b) Gaze and hand data.
(c) All data streams: gaze, head, and hands.
Figure 5: ROC curves showing the effect of patchwork privacy protections to different combinations of data streams on biometric verification.